OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 610! If four times the number of one-step equations minus twice the number of two-step equations is equal to 340, how many one-step equations auditioned today?

OpenStudy (zepp):

From the first sentence, you can say that One-Step Equations + Two Steps Equations = 610 If 4 * One Step - 2 * Two Steps Equation = 340 So lets the # of One step equation be x and # of two steps equation be y x - 7 = 610 4x - 2y = 340 And by using the substitution method, you can find the solution ;)

OpenStudy (anonymous):

let's say one-step equations is x and two-step equations is y You should have two equations 1) x+y=610 2)4x-2y=340 Solve the first for x and plug that into equation 2 to find y.

OpenStudy (zepp):

x + y = 610 4x - 2y = 340 It would be x + y = 610 x = 610 - y Substitute x into the second equation 4(610 - y) - 2y = 340 2440 - 4y - 2y = 340 2440 - 6y = 340 2440 - 340 = 6y 2100 = 6y y = 2100/6 y = 350 Plug it back it this first equation x + y = 610 x + 350 = 610 x = 260

OpenStudy (anonymous):

thank you.

OpenStudy (zepp):

x + y = 610 4x - 2y = 340 Then verify you answer, x= 260 y = 350 x + y = 610? 260 + 250 = 610 4(260) - 2(350) = 340? 340 = 340 Np ;)